# SURE-VQA: Systematic Understanding of Robustness Evaluation in Medical VQA Tasks

### v1

@@ 3.2 Human Rater Study
To ensure that the scores assigned by Mistral align with human judgment, we conduct a human rater study. For each dataset, we randomly selected 50 open-ended questions where the prediction did not exactly match the ground truth, as exact matches would automatically score highest by the hybrid metric (R2). Five human raters evaluated the questions, and we calculate the correlation between humans and Mistral’s scores using Kendall’s Tau (Kendall (1945)). Additionally, we report the correlation between humans and traditional metrics, and the inter-rater variability.
@@ Results
The results of the human rater study are presented in Figure 3, with detailed results available in Appendix B.1. Mistral demonstrates the highest correlation with human ratings on the SLAKE and OVQA datasets, with differences between Mistral and the best traditional metrics being $0.02$ and $0.13$, respectively. On the MIMIC dataset, traditional metrics perform slightly better, with the difference between Mistral and traditional metrics ranging from $0.08-0.18$. This variation can be attributed to several factors. In the OVQA dataset, for example, the structure of the questions and answers makes traditional metrics more prone to failure. Tokens like ”fracture,” ”left”/”right”, or specific bone names often match between the ground truth and predictions, despite significant differences in other tokens. An example of such a mismatch is illustrated in 4(a). On the other hand, MIMIC answers are highly structured, particularly for open-ended questions, where a fixed set of classes is listed in a comma-separated format. In these cases, traditional metrics perform well because token matching tends to be more accurate especially when the classes have distinct tokens.
Despite these limitations, Mistral’s performance on the MIMIC dataset remains strong, and its correlation with human ratings is in a similar range to that observed for the SLAKE dataset. Moreover, the failures of Mistral on MIMIC are not complete failures, as shown in 4(b). Here, Mistral does not put out the opposite than it should but rather a wrong tendency. Further, in the first example shown on the left, also some of the traditional metrics fail. In summary, although there are instances where traditional metrics show higher correlations with human scores, Mistral proves to be generally more robust and less prone to complete failures. Its correlation remains consistently high across all datasets, confirming its suitability as a reliable metric for evaluation in VQA tasks.
@@
Figure 3: Results of the Human Rater Study. Human interrater correlation is calculated between five human raters. We use Kendall’s Tau (Kendall (1945)) for calculating the correlation.
(a) Failures of traditional metrics on the OVQA dataset
(b) Failures of Mistral on the MIMIC dataset
Figure 4: Qualitative Results of the Human Rater Study. For each sample, the question, ground truth, prediction, and expected score by human ratings are shown. On the bottom, for each automated metric, the absolute value is shown with an indication if it is high or low in the metrics range and an indication if the expectations from the human ratings are met.
4 Empirical Study on the Robustness of Fine-Tuning Methods
@@
To show the relevance of SURE-VQA we performed an empirical study comparing the robustness of various FT methods under realistic shift in medical VQA. This is especially important since practitioners should be informed about the differences of the FT methods not only in terms of their performance but also how robust they are when selecting a method.
4.1 Study Design
We utilize (image, text) datasets from the medical domain, splitting them so that the training and testing distributions differ. As our base model, we employ LLaVA-Med 1.5, a state-of-the-art medical VLM (Li et al. (2023)). We fine-tune the model using four methods: full FT, prompt tuning (Lester et al. (2021)), LoRA (Hu et al. (2021)), and $\mathtt{(IA)^{3}}$(Liu et al. (2022)). Hyperparameters for the PEFT methods are selected based on the full training set and corresponding validation set for each dataset. Details regarding the hyperparameter search can be found in Appendix C.1. To measure robustness, we split the data into i.i.d. training and i.i.d. and OoD test sets, as outlined in section 3.1, thereby fulfilling R1. We then evaluate the performance of the VLM using Mistral as an evaluator, fulfilling R2. For robustness measurement, we calculate the relative robustness (RR) (Chen et al. (2023)), defined as $RR=1-\Delta P/P_{I}$, where $\Delta P=(P_{I}-P_{O})$, and $P_{I}$ is the i.i.d test performance and $P_{O}$ is the OoD test performance. For a better interpretation of the results, we compare them against relevant sanity baselines as described in R3.
4.2 Results
The results of the FT robustness study can be seen in Figure 5 and Figure 6. Detailed tables with the results are shown in Appendix C.4.
Figure 5: Results of the FT Robustness Study on the i.i.d. and OoD Test Set. Reported results show the mean over three seeds (exception: no FT, full FT) with the standard deviation for the non-baselines. Mistral Accuracy refers to the accuracy being rated by Mistral, meaning it assigns 00 or $1$ to the output. For MIMIC, the most frequent sanity baseline can not be calculated as too few questions match the training set.
Figure 6: Results of the FT Robustness Study Focusing on the Relative Robustness (RR). Since the study focuses on comparing the PEFT method and our definition of OoD holds for these methods, the no FT baseline is excluded. (a) RR on the three datasets for all FT methods. Results show the mean and standard deviation over three seeds (exception: full FT). (b) RR Ranking of the methods. (c) Standard deviation between shifts vs. standard deviation between FT methods.
Comparison to Sanity Baselines.
@@ 4.2 Results
Generally, the PEFT models outperform the no fine-tuned models, the most frequent baseline, and their respective no image baselines on the i.i.d. datasets. This is in contrast to full FT, which mostly does not outperform the most frequent baseline. However, the gap between the no image baselines and the PEFT models varies across datasets. For example, it is largest on the SLAKE dataset, averaging 24%, but much smaller on the OVQA dataset at around 11%, and similarly small on the MIMIC dataset at 7%. Interestingly, the MIMIC dataset yields relatively poor results with fine-tuned models, averaging just 61.2% mistral accuracy on closed-ended questions and 3.2 Mistral score on open-ended ones on the i.i.d. set. Specifically, for prompt tuning, the performance of the model fine-tuned with images is almost identical to its no image counterpart, showing only about a 2% improvement. This indicates that, for the MIMIC dataset, the image encodings contribute little to the question-answering task, suggesting that the vision encoder is not able to extract meaningful information from the images. This could be because the MIMIC dataset focuses on detailed chest X-ray images and the image encoder does not have this fine-grained expertise in such a specific task. While the no image baseline and the most frequent baseline both mostly perform better than random classifier, this effect is particularly noticeable for the closed-ended questions on OVQA, where the no image baseline achieves on average 72%, and the most frequent baseline 73.6%. This suggests that the model can often rely on learned question-answer correlations without needing the images.
Comparison Between Datasets.
For closed-ended questions, the average i.i.d. performance of the PEFT models is similar between the SLAKE (86%) and OVQA (83.5%) datasets. However, the average gap between these models and the no image sanity baseline is larger on SLAKE compared to OVQA as mentioned above. This suggests that image information has a greater impact on overall model performance in the SLAKE dataset. A reason for this could be that the ratio of unique questions is smaller in the OVQA dataset (see Table 13, Appendix C.3.4), and also the most frequent baseline is performing better on this dataset. This suggests that there are more repetitive questions and higher bias, making the model prone to shortcut learning. Continuing on closed-ended questions, the PEFT models outperform the no-finetuned model by a larger margin on OVQA, with a 41% improvement on average, compared to 29% on SLAKE. This suggests that fine-tuning has a greater impact on the OVQA dataset. For open-ended questions, the average i.i.d. performance is the best on the SLAKE dataset with a Mistal score of 4.22. In contrast, the performance on the MIMIC dataset seems generally insufficient for practical use. For open-ended questions, the average performance improvement over the no image baseline across all PEFT methods is just 9%, and for closed-ended questions, it’s only 5%. Especially for prompt tuning, there is a 3% decrease for closed-ended ones.
Comparison between Full FT and PEFT Methods.
In our evaluation, full FT demonstrated inferior performance compared to PEFT methods on the SLAKE and OVQA datasets and only partially matched or slightly surpassed LoRA on some splits of the MIMIC dataset. Notably, the performance of full FT generally improved with increasing dataset size, performing worst on SLAKE, slightly better on OVQA, and best on MIMIC, aligning with observations by Dutt et al. (2024). Further, we observed a significant failure of full FT robustness for the OVQA question type shift. Overall, as full FT neither outperforms PEFT methods in terms of performance nor robustness, we agree with Dutt et al. (2024) that PEFT methods are particularly well-suited for medical and low-data scenarios.
Comparison Between PEFT Methods.
@@
On the i.i.d. set, LoRA is consistently the best PEFT method, achieving an average accuracy of 81.8% on closed-ended and a Mistral score of 3.6 on open-ended questions across different datasets and shifts. However, the performance differences between LoRA and other PEFT methods are quite small, especially considering that these other methods require fewer parameters to train. $\mathtt{(IA)^{3}}$performs an average of 3.5% below LoRA, while prompt tuning shows a 5% lower performance on average across the datasets. Overall, robustness within methods is more homogeneous, while there is more variance across dataset shifts. This suggests that the type of shift has a greater impact on robustness than the choice of the fine-tuning method as visualized in Figure 6c). Only on the OVQA dataset, the inter-method variability is higher than the inter-shift variability, but this is due to the failure in robustness of full FT. Only comparing the PEFT methods, also for OVQA the inter-shift variability is lower, shown in Appendix C.4, Figure 12. Besides that, none of the fine-tuning methods consistently outperforms the others in terms of robustness as seen in figure Figure 6b), where the rank of each method is depicted with respect to their RR. As one exceptional outlier of the PEFT methods, on closed-ended questions in SLAKE, LoRA demonstrates significantly lower robustness compared to other methods, with an RR of 48%, compared to 65% for prompt tuning and 68% for $\mathtt{(IA)^{3}}$.
@@
The robustness trends vary between datasets, as shown in Figure 6a. On SLAKE, models demonstrate greater robustness on open-ended questions compared to closed-ended ones, while the opposite is true for OVQA, where models perform more robustly on closed-ended questions. This pattern is especially clear in question type shifts, where RR on SLAKE is 54% for closed-ended questions and increases to 101% for open-ended. In contrast, on OVQA, RR is 94% for closed-ended and drops to 55% for open-ended questions. At the same time, the question type shift seems most severe if the model performance drops. The differing behavior in SLAKE and OVQA datasets arises from the alignment of OoD questions with training data. While the OoD questions are not included in the training set, training questions help the model capture the necessary information to address them. This alignment is observed in SLAKE for open-ended questions and in OVQA for closed-ended ones. The population shifts on MIMIC did not affect the models’ robustness, i.e. the models seem robust against such shifts since they show over 100% RR. However, since the performance on the MIMIC dataset is generally insufficient, it is questionable if this observation would hold on higher i.i.d. performance. Future experiments could investigate whether the low performance or the kind of shift is causing this behavior.
@@ 5 Conclusion and Take-aways
While in section 2 we derived R1-R3 from flaws in the currrent literature, our study provides empirical evidence for their importance. R1: In an ablation study (Appendix D), we show that corruption shifts do not necessarily translate to real-world shifts, thereby justifying our claim to also work with more real-world shifts. R2: We present several critical failures of traditional token-matching metrics and prove the applicability of our LLM evaluation setup by a human rater study. R3: We show that some sanity baselines that do not use the image information already perform surprisingly well. This highlights two aspects: 1) As stated in R3, reporting such sanity baselines is crucial for understanding the true multi-modal performance of a VLM, beyond its language-only capabilities. 2) This observation further suggests that we need more elaborate data sets and tasks in medical VQA that minimize the potential for shortcut learning based solely on the language content. Achieving this may involve incorporating greater linguistic variability in questions, as seen in Bae et al. (2023), where questions were rephrased using GPT-4, and ensuring a broad range of semantic differences in the questions. Progress can be tracked as a low performance of the no-image sanity baseline.
@@
SURE-VQA, serves as a starting point for a comprehensive evaluation of robustness and it can be flexibly extended to new datasets, methods, and domains. Additionally, SURE-VQA can support method development aimed at enhancing the robustness of VLMs. In our study, we define OoD as a data shift w.r.t the FT data. However, our framework also allows to compare VLMs in a zero-shot setting without any FT, when simply re-defining OoD as a data shift w.r.t the pre-training data. However, such a definition becomes increasingly challenging to validate with foundation models, since the exact training data used is often not known. Notably, R2, and R3 go beyond robustness analysis and should be integral to any well-designed VLM study.
@@
Our exemplary study which compares the robustness of various FT methods reveals several key insights. While we confirm LoRA as the best-performing FT method on the i.i.d. dataset, no single FT method consistently outperforms the others in terms of robustness. Further, in line with findings of Dutt et al. (2024), we find that PEFT methods are more efficient than full FT in the lower data regimes, especially present in the medical domain. As another finding, robustness trends appear to be more consistent within FT methods than across different dataset shifts, indicating that the type of shift has a greater impact on robustness than the choice of FT method. This suggests that robustness alone is not a decisive factor when choosing a FT method. However, the type of data shift anticipated in the test set is crucial, as different shifts may uniquely challenge model performance. Additionally, the models generally exhibit robustness against population shifts. However, further investigation is needed to determine whether this is due to already low i.i.d. performance or only because of the nature of the shift.
@@
As mentioned above, SURE-VQA is a starting point and future work can include the investigation of more datasets, shifts, methods, and models. A key research direction is the development of additional VQA datasets, particularly in the medical domain. This would not only address the need for greater diversity in question types but also improve the clinical relevance of the questions posed. While current datasets cover various question types, questions about scanning modalities, for example, may be less valuable to clinicians. More relevant questions might include for example questions related to prognosis, such as the potential spread of a tumor. We believe that collaboration with clinicians could help define and incorporate a broader set of clinically meaningful questions into future datasets. Additionally, our framework provides valuable insights into the factors affecting the robustness of VLMs. Future work could explore methods for enhancing robustness, like in Yoon et al. (2024); Ma et al. (2024). Finally, the underperformance of LLaVA-Med, one of the state-of-the-art models in medical VQA, on the MIMIC-CXR-VQA dataset indicates that there is significant room for improvement in medical VLM development. Recent work by Chen et al. (2024b) has focused on building foundation models specifically for chest X-ray data, using datasets like MIMIC-CXR-VQA. However, future efforts could aim to develop more robust models capable of handling multiple modalities across a wider range of clinical scenarios.

### v2

@@ 3.3.1 R1: Comparison between Realistic Shifts and Corruption Shifts
To compare the effect of realistic shifts (R1) against artificial shifts (corruption shifts), we perform a robustness study on the SLAKE dataset, using a simplified setup that involves only the LLaVA-Med model fine-tuned on SLAKE with $\mathtt{(IA)^{3}}$. This study aims to support the claims outlined in (R1) within the context of our dataset setup. As illustrated in Figure 20, the model is first fine-tuned on an i.i.d. training set. Then, we test it on the i.i.d. test set and two different OoD test sets—one reflecting realistic shifts, and the other containing artificially corrupted images. For realistic shifts, we use the original i.i.d. and OoD splits defined in section 3.1. For artificial (corruption) shifts, the OoD test set is generated by applying image corruptions like blur, brightness, and noise at varying severity levels to the i.i.d. test images. Each corruption is applied with a probability of 0.5, and at least one corruption is applied per image.
@@ Results
As seen in Figure 3, the realistic modality and question type shifts exhibit lower relative robustness compared to artificial shifts (medium strength), meaning artificial shifts fail to accurately capture the challenges presented by realistic shifts. This finding highlights the importance of evaluating robustness under realistic conditions, as adopted in our study. While corruption-based shifts may offer a valuable complementary perspective, these results emphasize that they should not be considered the only source for evaluating robustness. Further details and experiments on other corruption levels are provided in Appendix D.
Figure 3: Comparison between Realistic shifts and Corruption shifts on SLAKE Dataset Finetuned on $\mathtt{(IA)^{3}}$. For realistic shifts, we use the same i.i.d. and OoD training and test sets as described in Section 3.1. For corruption shifts, the i.i.d. training and test sets remain the same as in the realistic setup. The only difference is the OoD test set, which is created by adding one or more corruptions such as blur, brightness changes, or noise to the i.i.d. test images at varying levels of severity.
3.3.2 R2: Human Rater Study
Study Design
To ensure that the scores assigned by the LLM align with human judgment, we conduct a human rater study. We use this to compare various LLMs for rating and select the one that aligns best with human judgment. For each dataset, we randomly selected 100 open-ended questions where the prediction did not exactly match the ground truth, as exact matches would automatically score highest by the hybrid metric (R2). Five human raters evaluate the questions, and we calculate the correlation between humans and the LLM using Kendall’s Tau (Kendall (1945)). Additionally, we report the correlation between humans and traditional metrics, and the inter-rater variability.
Results
@@
Figure 4: Results of the Human Rater Study. Human interrater correlation is calculated between five raters. We use Kendall’s Tau Kendall (1945) as correlation measure.
Among the LLMs evaluated, Gemma demonstrates the highest correlation with human ratings, as detailed in Figure 10 (Appendix B.1). Thus, Gemma is selected as the evaluator, and we present the results of the human rater study using Gemma in the following section. The correlations are shown in Figure 4, with detailed results available in Appendix B.3. Gemma demonstrates the highest correlation with human ratings on the SLAKE and OVQA datasets, with differences between Gemma and the best traditional metrics being $0.15$ and $0.21$, respectively. Interestingly, on these datasets, the human interrater correlation is not the highest compared to correlation with other metrics. This suggests that the metrics fall between the evaluations of the human raters, resulting in a stronger correlation with the average human rating than between the individual raters themselves. On the MIMIC dataset, traditional metrics perform slightly better, with the performance gap between Gemma and the traditional metrics ranging from $0.03$ to $0.08$. This variation can be attributed to several factors. In the OVQA dataset, for example, the structure of the questions and answers makes traditional metrics more prone to failure. Tokens like "fracture," "left"/"right", or specific bone names often match between the ground truth and predictions, despite significant differences in other tokens (see 11(a) for an example). On the other hand, MIMIC answers are highly structured, particularly for open-ended questions, where a fixed set of classes is listed in a comma-separated format. In these cases, traditional metrics perform well because token matching tends to be more accurate especially when the classes have distinct tokens.
Despite these limitations, Gemma’s performance on the MIMIC dataset remains strong, and its correlation with human ratings is even higher than on the other two datasets. Moreover, the failures of Gemma on MIMIC are not complete failures, as shown in 11(b). In summary, although there are instances where traditional metrics show higher correlations with human scores, Gemma proves to be generally more robust and less prone to complete failures. Its correlation remains consistently high across all datasets, confirming it as a reliable metric for evaluation in VQA tasks. However, even though Gemma shows high correlation with human judgments, it is important to note that LLMs may carry internal biases. These biases might come from underrepresentation of detailed medical knowledge, recent medical advancements, and diverse socio-demographic groups in LLMs training procedure and data. As a result, such factors can influence the model’s scoring behavior in evaluation tasks. That’s why it is crucial to conduct human-rater studies and choose an LLM that aligns closely with human judgment to minimize these effects especially when applied to a new use case.
3.3.3 R3: Performance of Sanity Baselines
Study Design
To evaluate the performance of sanity baselines, we compare a VLM not using the image information (no image sanity baseline) with a VLM using the image information. Specifically, we investigate how often the model using the images actually outperforms the model that is not using the images. This allows us to evaluate the model’s susceptibility to shortcut learning, where it may rely predominantly on language cues while neglecting visual information. While this section analyses the no image baseline in detail, the general confirmation of the other sanity baselines is part of our large-scale study in section 3.4.
Results
As seen in Figure 5, overall, the sanity baseline of not using the image information performs surprisingly well. Ideally, in the absence of shortcut learning, the model utilizing image information should almost consistently outperform its image-agnostic counterpart or at least tie on the individual questions. However, this is not the case on any of the datasets. While the win-to-loss ratio is overall positive for the model using the image information, the margin is narrower than anticipated. Most notably, on the MIMIC dataset, the model using the image information is even outperformed on up to 25% of the questions. While on the SLAKE dataset, the win-to-loss ratio is highest for the model using the image information, it is the worst on the MIMIC dataset, and on the OVQA dataset, the models tie the most. Further results on the other sanity baselines as well as potential reasons for performance differences between datasets are provided in section 3.4.
Figure 5: Comparison between the Model using the Images and the Model not using the Images. For each PEFT method, the answer scores of one model using the images are compared to the corresponding answers of the model not using the images. Lose = Model using the images loses, Tie = Both answers are equally good, Win = Model using the images wins. Closed i.i.d. / OoD = Closed-Ended questions, Open i.i.d. / OoD= Open-Ended questions. Figure layout inspired by Jeong et al. (2024)
3.4 Empirical Study on the Robustness of Fine-Tuning Methods
@@
To show the relevance of SURE-VQA we perform an empirical study comparing the robustness of various FT methods under realistic shifts in medical VQA. This is important since practitioners should be informed about the differences between the methods not only in terms of their performance but also how robust they are when selecting a method. The results of the FT robustness study can be seen in Figure 6, Figure 7 and Figure 8. Detailed tables with the results are shown in Appendix C.5. Further, an ablation on the impact of multimodal shifts is shown in Appendix F.
Comparison between Medical and Non-Medical Base Models.
Figure 6 shows the comparison between the medical (LLaVA-Med 1.5) and the non-medical base model (LLaVA 1.6), aggregated over all FT methods. The results indicate that medical and non-medical models often tie in their answers, with ties ranging from 64% to 90%. On SLAKE and OVQA, the medical model generally achieves a slightly higher number of wins. On the MIMIC dataset, the medical model is slightly outperformed on closed-ended questions but shows marginally better performance on open-ended ones. However, the differences in wins and losses between the two models are mostly minor. This suggests that there is no clear winner between medical and non-medical base models when fine-tuning. Similar results have been shown by Jeong et al. (2024) in zero- and few-shot settings, however, in their experiments, the medical model performed slightly worse. To make the discussion about the following results easier to parse, we detail the results of the medical base model in the main paper and show detailed results on the non-medical base model in Appendix C.4.
Comparison Between Datasets Considering the Sanity Baselines
@@
Generally, the PEFT models outperform the no fine-tuned models, the most frequent baseline, and their respective no image baselines on the i.i.d. datasets. This is in contrast to full FT, which often does not outperform the most frequent baseline. Notably, while the no image baseline and the most frequent baseline both mostly perform better than random classifier, this effect is particularly noticeable for the closed-ended questions on OVQA, where the no image baseline achieves on average 72%, and the most frequent baseline 73.6%. Consequently, the gap between the no image baselines and the PEFT models is smaller for OVQA (18%) compared to SLAKE (33%). This indicates that the model can often rely on learned question-answer correlations without needing the images, likely because the OVQA dataset has the least unique questions (see Table 4, Appendix C.1.4). For the MIMIC dataset, the gap between the no image baseline and the PEFT methods is even smaller than for OVQA, with only an improvement of 11% for the PEFT methods over the baselines. However, as Table 4 suggests, MIMIC does not have many repetitive questions. Thus, the small gap between PEFT models and the no image baseline might mean that image encodings contribute minimally to the VQA task. This is likely because the MIMIC dataset focuses on detailed chest X-ray images, and the image encoder may lack this fine-grained expertise. Regarding the no-finetuned model baseline, PEFT models outperform the no-finetuned model on the closed-ended questions by the largest margin on OVQA, with a 43% improvement on average. However, this trend differs on the open-ended questions, where the difference between the fine-tuned models and the no-finetuned model seems larger on SLAKE and MIMIC. To summarize, the models overall seem to learn most on the SLAKE dataset, which also shows the highest i.i.d. performance on the PEFT methods with an average accuracy of 86.2% and a Gemma score of 4.3. In contrast, the closed-ended performance on the MIMIC dataset seems generally insufficient for practical use (61%).
Figure 6: Comparison between the Medical and the Non-Medical Base Model. For each FT method, the answer scores of one medical experiment are compared to the corresponding answers of the non-medical experiment. Lose = Medical model loses, Tie = Both answers are equally good, Win = Medical model wins. Closed i.i.d. / OoD = Closed-Ended questions, Open i.i.d. / OoD= Open-Ended questions. Figure layout inspired by Jeong et al. (2024).
Figure 7: Results of the FT Robustness Study on the i.i.d. and OoD Test Set. Reported results show the mean over three seeds (exception: no FT, full FT) with the standard deviation for the non-baselines. Gemma Accuracy refers to the accuracy being rated by Gemma. For MIMIC, the most frequent sanity baseline can not be calculated as too few questions match the training set.
Figure 8: Results of the FT Robustness Study Focusing on the Relative Robustness (RR). Since the study focuses on comparing the FT methods and our definition of OoD holds for these methods, the no FT baseline is excluded. (a) RR on the three datasets for the FT methods. Results show the mean and standard deviation over three seeds (exception: full FT). (b) RR Ranking of all FT methods. (c) Standard deviation between shifts vs. standard deviation between FT methods.
Comparison Between FT Methods.
@@
When comparing full FT with PEFT methods, we see that full FT neither outperforms PEFT methods in terms of performance nor robustness, aligning with the observations of Dutt et al. (2024) that PEFT methods are particularly well-suited for medical and low-data scenarios. From the PEFT methods, on the i.i.d. set, LoRA is consistently the best PEFT method, achieving an average accuracy of 80.6% on closed-ended and a Gemma score of 3.6 on open-ended questions across different datasets and shifts. However, the performance differences between LoRA and other PEFT methods are quite small, especially considering that these other methods require fewer parameters to train. Overall, robustness within methods is more homogeneous, while there is more variance across dataset shifts. This suggests that the type of shift has a greater impact on robustness than the choice of the fine-tuning method as visualized in Figure 8c). Only on the OVQA dataset, the inter-method variability is higher than the inter-shift variability, but this is due to the failure in robustness of full FT. Only comparing the PEFT methods, also for OVQA the inter-shift variability is lower, shown in Appendix C.5, Figure 18. Besides that, none of the fine-tuning methods consistently outperforms the others in terms of robustness as seen in Figure 8b), where the rank of each method is depicted with respect to their RR. As one exceptional outlier of the PEFT methods, on closed-ended questions in SLAKE, LoRA demonstrates significantly lower robustness compared to other methods, with an RR of 48%, compared to 65% for prompt tuning and 68% for $\mathtt{(IA)^{3}}$.
@@
The robustness trends vary between datasets, as shown in Figure 8a). On SLAKE, models demonstrate greater robustness on open-ended questions compared to closed-ended ones, while the opposite is true for OVQA. This behavior results from the related question patterns between OoD questions and the training data. Although the OoD questions are not part of the training set, similar training questions offer enough context for generalization, seen in SLAKE for open-ended questions and in OVQA for closed-ended ones. Similar analysis on a conducted ablation study is provided in Appendix E. This pattern is especially clear in question type shifts, where RR on SLAKE is 54% for closed-ended questions and increases to 98% for open-ended. In contrast, on OVQA, RR is 94% for closed-ended questions and drops to 63% for open-ended ones. This means that the question type shift seems most severe if the model performance drops. The population shifts on MIMIC did not affect the models’ robustness, i.e. the models seem robust against such shifts since they show over 100% RR. However, since the performance on the MIMIC dataset is generally insufficient, it is questionable if this observation would hold on higher i.i.d. performance. Future experiments could investigate whether the low performance or the kind of shift is causing this behavior.
@@ 4 Conclusion and Take-aways
While in section 2 we derived R1-R3 from flaws in the current literature, our study provides empirical evidence for their importance. R1: In a study (section 3.3.1 and Appendix D), we show that corruption shifts do not necessarily translate to real-world shifts, thereby justifying our claim to also work with more real-world shifts. R2: We present several critical failures of traditional token-matching metrics and prove the applicability of our LLM evaluation setup by a human rater study in section 3.3.2. R3: We show that some sanity baselines that do not use the image information already perform surprisingly well (section 3.3.3 and section 3.4). This highlights two aspects: 1) As stated in R3, reporting such sanity baselines is crucial for understanding the true multi-modal performance of a VLM, beyond its language-only capabilities. 2) This observation further suggests that we need more elaborate datasets and tasks in medical VQA that minimize the potential for shortcut learning based solely on the language content. Achieving this may involve incorporating greater linguistic variability in questions, as seen in Bae et al. (2023), where questions were rephrased using GPT-4, and ensuring a broad range of semantic differences in the questions. Progress can be tracked as a low performance of the no-image sanity baseline.
@@
SURE-VQA serves as a starting point for a comprehensive evaluation of robustness and it can be flexibly extended to new datasets, methods, and domains. Additionally, SURE-VQA can support method development aimed at enhancing the robustness of VLMs. In our study, we define OoD as a data shift w.r.t the FT data. However, our framework also allows to compare VLMs in a zero-shot setting without any FT, when simply redefining OoD as a data shift w.r.t the pre-training data. However, such a definition becomes increasingly challenging to validate with foundation models, since the exact training data used is often not known. Notably, R2, and R3 go beyond robustness analysis and should be integral to any well-designed VLM study.
@@
Our exemplary study which compares the robustness of various FT methods reveals several key insights. With regard to robustness, we find that no single FT method consistently outperforms the others in terms of robustness. Furthermore, robustness trends appear to be more consistent within FT methods than across different dataset shifts, indicating that the type of shift has a greater impact on robustness than the choice of FT method. This suggests that robustness alone is not a decisive factor when choosing a FT method. Additionally, the models are generally robust to population shifts. However, further investigation is needed to determine whether this is due to low i.i.d. performance or because of the nature of the shift. Besides these insights into robustness, our comparison of medical and non-medical base models shows that both perform similarly when fine-tuned. Finally, in line with findings of Dutt et al. (2024), we find that PEFT methods are more efficient than full FT. Specifically, we confirm LoRA as the best-performing FT method on the i.i.d. dataset.
@@
As mentioned above, SURE-VQA is a starting point serving as a foundational framework and is not yet intended as a clinically deployable benchmark for medical VQA systems. Instead, it functions as a research tool to inform and guide future developments in the field. Future work can include the investigation of more datasets, shifts, methods, and models. Particularly, the exemplary study in this paper mainly focuses on radiological VQA datasets. Other medical image types like histopathology, surgical, or dermatology images could be investigated in future work as they may have different characteristics and thus may reveal other findings. Further, while this paper focuses on simple, computational efficient baselines that omit image input, future work could use Shapley values to gain additional insights into the role of images in the VQA task (Parcalabescu & Frank (2023)). Another key research direction is the development of additional VQA datasets, particularly in the medical domain. This would cover two needs: 1) The need for more diverse question types, mitigating shortcut learning during fine-tuning based on the language, and improving the generalization on unseen data and 2) improve the clinical relevance. While current datasets cover various question types, they might offer limited value to clinicians. For example, clinicians might find questions on prognosis and tumor spread more beneficial than scanning modalities. Therefore, collaborating with clinicians can help incorporate more clinically meaningful questions. Finally, the underperformance of two state-of-the-art VLMs, LLaVA and LLaVA-Med on the MIMIC-CXR-VQA dataset underscores the critical need for advancing the development of medical VLMs, particularly given that LLaVA-Med is specifically designed for the medical domain. Even though recent work by Chen et al. (2024b) has focused on developing foundation models specifically for chest X-ray data, there is still a need to develop more robust models capable of handling multiple modalities across a wider range of clinical scenarios.
Acknowledgements
This work was partly funded by Helmholtz Imaging (HI), a platform of the Helmholtz Incubator on Information and Data Science. We thank Dasha Trofimova for her help in the human rater study as well as insightful discussions and feedback.